Tickets for a play at the community theater cost $20 for an adult and $8 for a child. If 120 tickets were sold and the total receipts were $2040, how many of each type of ticket were sold?
Answer by ptaylor(2198)

Let x=number of adult tickets sold
Then 120-x=number of child tickets sold
Now we are told that the receipts for adult tickets ($20x) plus the receipts for child tickets ($8(120-x)) equals $2040. So our equation to solve is:
20x+8(120-x)=2040 get rid of parens
20x+960-8x=2040 subtract 960 from both sides
20x+960-960-8x=2040-960 collect like terms
12x=1080 divide both sides by 12
x=90-----------------------------number of adult tickets sold
120-x=120-90=30--------------------number of child tickets sold
ck
20*90+8*30=2040
1800+240=2040
2040=2040
also 90+30=120
120=120
Hope this helps-----ptaylor